OpenStudy (anonymous):

What is the standard deviation of the data set given below? 2, 3, 6, 9, 10 A. 10 squared B. 3.5 squared C. 12.5 D. 10 E. 3.5 F. 12.5 squared

OpenStudy (anonymous):

first find the mean

OpenStudy (anonymous):

the mean is 6

OpenStudy (anonymous):

actually lets divide by n-1, since you have sample data

OpenStudy (anonymous):

$$ \large st.dev = \sqrt{\frac{ ( 2- 6 )^2 + (3 - 6 )^2 + (6-6 )^2 +(9-6 )^2 }{5-1}}$$

OpenStudy (anonymous):

i got 13.8

OpenStudy (anonymous):

actually i got 3.4

OpenStudy (anonymous):

$$ \large{ st.dev = \sqrt{\frac{ ( 2- 6 )^2 + (3 - 6 )^2 + (6-6 )^2 +(9-6 )^2 }{5-1}} \\ ~~~~~~~~~~= \sqrt{\frac{50}{4} } }$$

OpenStudy (anonymous):

I got 3.53

OpenStudy (anonymous):

squared?

OpenStudy (anonymous):

we are taking the square root, so it isn't squared